OpenStudy (xapproachesinfinity):

let c and s both be continuous function if \[s'(x)=c(x)~~,~~ c'(x)=-s(x) ~~and~~ s(0)=0,~~c(0)=1\] prove that \[[s(x)]^2+[c(x)]^2=1\]

OpenStudy (xapproachesinfinity):

i know this sin and cos i'm stuck oh how to show they are

OpenStudy (xapproachesinfinity):

the conditions satisfy sin and cos functions

OpenStudy (xapproachesinfinity):

i was thinking of adding the two derivative

OpenStudy (anonymous):

\[\begin{cases}s'(x)=c(x)&(1)\\c'(x)=-s(x)&(2)\end{cases}\] Differentiating (1) gives \(s''(x)=c'(x)\). Substituting into (2) gives \[s''(x)=-s(x)~~\iff~~s''(x)+s(x)=0\] which is a second order ODE with characteristic equation \(r^2+1=0\), roots \(r=\pm i\), which generate the fundamental solution set \(\{\cos x,\sin x\}\), like you said. Alternatively, you can solve this as you would a system of ODEs. \[\begin{pmatrix}s(x)\\c(x)\end{pmatrix}'=\begin{pmatrix}0&1\\1&0\end{pmatrix}\begin{pmatrix}s(x)\\c(x)\end{pmatrix}\] The coefficient matrix has eigenvalues \(\lambda=\pm i\), which give eigenvectors \(\vec{\eta}=\begin{pmatrix}\mp i\\1\end{pmatrix}\), again giving you the same solution set.

OpenStudy (anonymous):

Slight mistake, the bottom left 1 in the coefficient matrix should be negative.

OpenStudy (xapproachesinfinity):

i got what you did but how would we prove the identity (this is not a DE question) it is cal2 problem

OpenStudy (xapproachesinfinity):

you know i kind of went to that DE too but idk how it could help that way

OpenStudy (xapproachesinfinity):

my prof said do implicit differentiation to the identity we want to prove and some other stuff, but i disagree with him we need to prove that holds so we can use the identity itself we need to arrive at it

OpenStudy (xapproachesinfinity):

can't*

OpenStudy (anonymous):

Suppose we try a reverse approach to what your professor suggested. \[\begin{align*}\begin{cases}s'(x)=c(x)\\c'(x)=-s(x)\end{cases}~~\implies~~-s(x)s'(x)&=c(x)c'(x)\\ -2s(x)s'(x)&=2c(x)c'(x)\\ 0&=2c(x)c'(x)+2s(x)s'(x)\\ 0&=\frac{d}{dx}\left[(c(x))^2+(s(x))^2\right]\\ C&=(c(x))^2+(s(x))^2 \end{align*}\] Now when \(x=0\), you have \[\begin{align*}C&=(c(x))^2+(s(x))^2\\ C&=1^2+0^2\\ C&=1\end{align*}\]

OpenStudy (xapproachesinfinity):

yes that's good! that is what i was looking for i thought of summing both equations but i didn't go further

OpenStudy (xapproachesinfinity):

thanks

OpenStudy (anonymous):

yw

OpenStudy (xapproachesinfinity):

i was trying with a similar problem hehe \[f(x+y)=f(x)f(y)-g(x)g(y)~~and ~~g(x+y)=f(x)g(y)+g(x)f(y)\] and again i have to prove \[[f(x)]^2+[g(x)]^2=1\]

OpenStudy (xapproachesinfinity):

there are similar we have f(0)=1 g(0)=0

OpenStudy (xapproachesinfinity):

i will see how i would solve this one!

OpenStudy (anonymous):

I'm not too experienced with functional equations, but if I were to guess you might be able to find a form of a partial derivative if you play around with the terms.

OpenStudy (xapproachesinfinity):

hmm partial derivative! unfortunately i forgot how to play with calc3 stuff i will see what i can do :) thanks

OpenStudy (xapproachesinfinity):

my thought for now is to let x=y and then see what i can pull from that

OpenStudy (xapproachesinfinity):

trying this \[\begin{cases}f(x+y)=f(x)f(y)-g(x)g(y)\\g(x+y)=g(x)f(y)+g(y)f(x)\end{cases} \Longrightarrow \begin{cases}[f(x+y)]^2=[f(x)f(y)-g(x)g(y)]^2\\ [g(x+y)]^2=[g(x)f(y)+g(y)f(x)]^2\end{cases}\]

OpenStudy (xapproachesinfinity):

but this is not using differentiation lol

OpenStudy (xapproachesinfinity):

if i add them up i should get \[[f(x+y)]^2+[g(x+y)]^2=[f(x)+g(x)][f(y)+g(y)]\] since g(0)=0 and f(0)=1 then we 1

OpenStudy (xapproachesinfinity):

eh but this a flow didn't pay attention to left hand side lol

OpenStudy (xapproachesinfinity):

how about using differentiation from that point?

OpenStudy (xapproachesinfinity):

eh that must be incorrect if i let y=0 \[[g(x)]^2+[f(x)]^2=f(x)+g(x)\] this wouldn't make any sense darn

OpenStudy (xapproachesinfinity):

hello maki san :)

OpenStudy (xapproachesinfinity):

by the way the question we are doing is on the comment 14th

OpenStudy (xapproachesinfinity):

@Kainui have a look at this :)